OpenStudy (anonymous):

Jacob with his mower can mow and trim Mr. Clark's yard in 2 x 1/2 hours, while Nicole with her mower can mow and trim the same yard in just 2 hours. If they work together on this yard and keep out of each other's way, in how many minutes can they complete the yard?

OpenStudy (anonymous):

Can anybody please explain how to solve this?

OpenStudy (anonymous):

1 hour and 7 minutes

OpenStudy (anonymous):

can you please explain?

OpenStudy (jdoe0001):

1 hour 6 minutes 46 seconds to be exact :P

OpenStudy (jdoe0001):

jacob takes 2 hours and a half, that is 150mins nicole takes 2 hours only, that is 120mins how much does jacob do in 1 hr? that is in 60 secs? so what's 60secs of 150? 60/150 = 2/5 so jacob in 1 hr has done only 2/5 of the yardwork what about nicole? how much has she done in 1 hr? well, 60/120 = 1/2 so she's really done half the yard in 1hr how will they fare together? add their rate 2/5 + 1/2 and use the ratio of 1 hr over the total length of time, that is \(\bf \cfrac{2}{5}+\cfrac{1}{2}=\cfrac{1}{\textit{total time both will take}}\)

OpenStudy (jdoe0001):

then just solve for "total time both will take"

OpenStudy (mathstudent55):

They will work together for x time. If Jacob can do the job in 2.5 hours, he does x/2.5 of the job. If Nicole can do the job in 2 hours, she does x/2 of the job. x is the same for both because when they work together, they work the same amount of time. Adding their work together, they do the entire job, 1. \(\dfrac{1}{2.5}x + \dfrac{1}{2}x = 1 \) x = 1.1111... hours The time is \(66 \dfrac{2}{3} \) minutes.

OpenStudy (anonymous):

thanks a lot:)!